What am I supposed to do with this? Am I supposed to follow an order of operation?
I am having problems with this. Please help me.
Find all the numbers for which the rational expression is undefined:
a. (x+3)/[(x–2)(x+5)]
b. 7/(x²+81)

Answer by AnlytcPhil(1810)

What am I supposed to do with this? 
Am I supposed to follow an order of operation?
I am having problems with this. Please help me.
Find all the numbers for which the rational expression is undefined: 
a. (x+3)/[(x–2)(x+5)]

b. 7/(x²+81)

----------------------------------------------------

You're supposed to know that a fraction (rational expression) is undefined
if its denominator is equal to 0.

So you set the denominator = 0 and find out what values will make the
fraction undefined. You can ignore the numerator because it has no
effect on whether a fraction (rational expression) is defined or
undefined.

The denominator of (a.) is (x-2)(x+5)

So set (x-2)(x+5) equal to 0 and solve:

(x-2)(x+5) = 0

Set each factor = 0

x-2 = 0 gives x = 2
x+5 = 0 gives x = -5

So 2 and -5 are the values which cause the 
fraction to be undefined.

----------------------------------------

The denominator in part (b.) is x²+81

We set x²+81 = 0
          x² = -81

But if we take the square root of a negative
number we do not get a real number. So there
is no real solution to x²+81=0.  Therefore
(b.) is never undefined.  

Edwin